OpenStudy (anonymous):

Use trig substitution to integrate dx / x square root of 5 - x^2

OpenStudy (anonymous):

\[dx\over\ \sqrt{5-x^{2}}\] is that the expression to be integrated?

OpenStudy (anonymous):

if so try \[x = \sqrt{5}\sin \theta\]